# Shedding Light on Gravity: Black Hole Shadows and Lensing Signatures in Lorentz Gauge Theory

### v1

@@ VII Conclusion
Furthermore, our analysis of the black hole shadow revealed its explicit dependence on $A_{0}$, exhibiting a pronounced sensitivity that could be tested with high-precision observations such as those from the Event Horizon Telescope. The computed shadow radius provides robust criteria for distinguishing Lorentz gauge gravity from standard General Relativity through direct imaging. These findings demonstrate the potential of gravitational lensing and shadow imaging as powerful observational tools to probe modified gravity theories. Our results thus contribute significantly toward establishing precise observational tests that can discriminate between LGT and classical GR, paving the way for deeper insights into the fundamental nature of gravity and spacetime structure.

### v2

@@ VII Conclusion
Furthermore, our analysis of the black hole shadow revealed its explicit dependence on $A_{0}$, exhibiting a pronounced sensitivity that could be tested with high-precision observations such as those from the EHT. The computed shadow radius provides robust criteria for distinguishing Lorentz gauge gravity from standard General Relativity through direct imaging. These findings demonstrate the potential of gravitational lensing and shadow imaging as powerful observational tools to probe modified gravity theories. Our results thus contribute significantly toward establishing precise observational tests that can discriminate between LGT and classical GR, paving the way for deeper insights into the fundamental nature of gravity and spacetime structure.